For scaffolding system shown, assume P1-50 lbf and P2-20 lbf. Perform force and momentum equilibrium balances for each bar toUsing MATLAB PLease
Answer #1

P→

Bar 1:

\small \sum F_y = 0\Rightarrow T_A+T_B = P_1+T_C+T_D

\small \Rightarrow T_A+T_B - T_C-T_D = 50

Moment about A:

\small \sum M = 0 \Rightarrow 9*T_B-1* T_C-4*T_D -5*P_1 = 0

\small \Rightarrow 9T_B - T_C-4T_D = 5*50 = 250

Bar 2:

\small \sum F_y = 0\Rightarrow -P_2+T_C+T_D = 0

\small \Rightarrow T_C+T_D = 20

Moment about C:

\small \sum M = 0 \Rightarrow P_2*1-T_D*3 = 0

\small \Rightarrow 3T_D = 20

Equations are:

\small T_A+T_B - T_C-T_D = 50

\small 9T_B - T_C-4T_D = 5*50 = 250

\small T_C+T_D = 20

\small 3T_D = 20

Putting the four equation in matrix form:

\small \begin{bmatrix} 1 & 1 &-1 & -1\\ 0 & 9 & -1 &-4 \\ 0 &0 &1 &1 \\ 0 & 0 & 0 & 3 \end{bmatrix}*\begin{bmatrix} T_A\\ T_B\\ T_C\\ T_D \end{bmatrix}= \begin{bmatrix} 50\\ 250\\ 20\\ 20 \end{bmatrix}

Matlab:

>>A=[1 1-1-1:09-1-4;0 0 1 1;0 0 0 31: B = [50 250 20 20]; >> T = inv (A) *B T= 37.7778 32.2222 13.3333 6.6667